OpenStudy (anonymous):

A painter mixes gallons of paint in a large cylindrical bucket so that there will be no difference in color among individual gallons. If a gallon of paint has a volume of approx 4000cm^3, what is the maximum number of whole gallons of paint the can be poured into the bucket? the raidus is 40cm and the hieght is 60

OpenStudy (anonymous):

This may be helpful: Area of a circle = pi * radius ^2 Volume of cylinder = a bunch of circles stacked on top of one = anotherheight * pi * radius ^2

OpenStudy (anonymous):

find the volume of the bucket(pi*radius^2*height) and then divide it by the volume of the gallon(4000). u will get the no. of gallons of paint

OpenStudy (anonymous):

i tried that and still didn't get the right answer

OpenStudy (anonymous):

75 gallons..

OpenStudy (anonymous):

thats not one of the choices the choices are A. 16 B 14 C. 18 D 24

OpenStudy (anonymous):

(h*pi*r^2)/4000 =(60*3.14*40*40)/4000 =75.36 round off to 75

OpenStudy (anonymous):

please check the numbers in the question and units also

OpenStudy (anonymous):

the raidius is 40 n the height is 60

OpenStudy (anonymous):

4000cm^3 is the approx vlo

OpenStudy (anonymous):

according to me answer must be 75

OpenStudy (anonymous):

ok thanks